# Quantum Circuit Optimization by Graph Coloring

### v1

@@ Abstract
Depth optimization of a quantum circuit consisting of commuting operations is shown to be reducible to the vertex coloring problem in graph theory. The reduction immediately leads to an algorithm for circuit optimization of commuting gates utilizing any coloring solver. To examine its applicability, known quantum circuits from the literature are optimized.
@@ 1 Introduction
Consider a quantum circuit consisting of finitely many unitary operations in which all operations are pair-wise commuting. Examples include instantaneous quantum polynomial [1, 2], commuting local Hamiltonian [3, 4], or certain implementations of an oracle function in quantum algorithms [5, 6]. Such a circuit will be called a commuting circuit hereafter. Now, suppose a quantum machine given to us can handle more than one commuting operation at the same time if certain conditions imposed by the machine are met. In other words, some operations are parallelizable depending on the machine. This concept of parallelizability leads to the notion of depth (of a circuit).
In implementing a function by the logic circuit, multiple choices for the circuit construction are often allowed, and thus the most beneficial one is better chosen [7]. Given that the number of bits is fixed, a circuit with a shorter depth is often considered efficient as it means the parallelizability is fully exploited. For this reason, among others, optimizing the depth of a circuit is an active area of research [7, 8, 9, 10, 11].
This work uncovers the relationship between the depth optimization of commuting circuits and the problem in graph theory that is known to be NP-hard. The relation then introduces a circuit optimization strategy in a natural way. The results are summarized as follows:
@@
    The strategy is applied to optimize circuits for finite field multiplications and additions by quantum Fourier transform (QFT). The first application exhibits the improvement upon Maslov et al.’s design [5] and the second one gives tradeoff options for Draper’s quantum addition [6].
The paper is organized as follows. Preliminaries are given in Section 2. Main idea and proofs are given in Section 3. Applications are examined in Section 4, and some discussions are noted in Section 5.
@@ 2.2 Gate Ordering Problem
Definition 1 (Commuting operations)
@@
Definition 2 (Depth of a circuit)
@@
Definition 3 (Gate ordering)
@@ 2.3 Graph coloring
Definition 4 (Proper coloring)
@@
Definition 5 (Vertex coloring problem)
Given a graph $G=(V,E)$, a vertex coloring problem is to find a colored graph $(V,E,c)\in\mathbf{Col}(G)$ such that |$|\{c(v):v\in V\}|$ is minimized.
@@ 2.4 Related works
There exist works on quantum circuit optimization utilizing coloring in graph theory, by looking qubits as vertices. For example, edge coloring of a graph has been used to optimize the depth of Clifford circuits [22], or binary field inversion [23]. Such an approach is natural in the sense that once wires (qubits) are viewed as vertices and gates are considered as edges interconnecting the wires, the result of edge coloring tells which gates are parallelizable because two edges being connected to the same vertex implies they share the same qubit as the reference. On the other hand, to our best knowledge, this work is the first attempt to make use of a graph in which quantum gates are vertices.
The optimization approach developed in this paper uses an algorithm for graph coloring as the core routine. Any graph coloring algorithm can be used, for example including ones from previous works [24, 25, 26, 27], and it will soon be shown that the output of the graph coloring is directly related with the optimized circuit.
@@ 3 Reduction from gate ordering to graph coloring
The gate ordering problem is reduced to the graph coloring problem in this section. The opposite direction is also given in Appendix 0.A, showing that the two problems are equivalent. Since the graph coloring problem is in NP-hard [14], it implies optimizing a commuting circuit is also hard.
@@ 3.1.2 Colored graph to circuit
Case (b) is straightforward, thus we leave a brief comment on the smaller case (a). Given a colored graph, unconnected vertices correspond to parallelizable gates, but the colors on these vertices can be different. Therefore, despite the different colors on the vertices, the corresponding gates from the output may happen to be in the same layer.
@@ 3.2 Reduction
Recall that $G_{\cal V}$ is the graph drawn by Algorithm 1 for given ${\bm{v}}$.
Let $B_{\cal V}=\mathbf{Col}(G_{\cal V})$, which is the set of all colored graphs of $G_{\cal V}$.
Claim 7
@@
$m_{\rm color}(b)=\min_{b^{\prime}\in B_{\cal V}}\big{(}m_{\rm color}(b^{\prime})\big{)}\enspace.$ (1)
@@
$m_{\rm color}(g({\bm{a}}))=m_{\rm depth}({\bm{a}})<m_{\rm depth}(f(b))\leq m_{\rm color}(b)\enspace,$
which contradicts to Eq. (1). $\square$
@@
4:$(V,E)\leftarrow\mathsf{ConstructGraph}({\bm{v}})$triangleright $V=\{v_{1},\ldots,v_{n}\}$
5:$i\leftarrow 1$, $c_{1}\leftarrow 1$, $d\leftarrow 1$triangleright $d$ for depth
@@
17:return $(V,E,c)$triangleright $c(v_{i})=c_{i}$
Suppose a circuit ${\bm{v}}$ is an input to $\mathsf{ToColoredGraph}$. Recall that the vertices are gates. Algorithm 3 sequentially identifies the parallelizable gates (vertices), and those vertices are identically colored in Line 7. Whenever a new color is introduced in Line 10, $d$ is also increased by 1. Since $d$ is increased only when $v_{j}$ is non-parallelizable with any of $v_{i},\ldots,v_{j-1}$ (deducible from Line 6), the final value of $d$ at the end is equal to the depth of the circuit. Therefore, we see that $\forall{\bm{a}}\in A_{\cal V}$, $m_{\rm depth}({\bm{a}})=m_{\rm color}(\mathsf{ToColoredGraph}({\bm{a}}))$. Time complexity of Algorithm 3 is $O(n^{2})$.
We have shown that Claim 7 does hold, thereby completing the reduction from gate ordering to graph coloring. A reduction for the opposite direction, from graph coloring to gate ordering, is less important as the main purpose of this work is to suggest a strategy for circuit optimization. Therefore we leave the opposite reduction in Appendix 0.A.
@@ 3.3 Circuit optimization strategy
It can be deduced from inspecting the reduction steps that an approximate solution of graph coloring leads to at least as good approximations to the gate ordering problem. Let ${\bm{v}}$ be a commuting circuit, ${\bm{v}}_{\rm opt}$ be the optimal circuit of ${\bm{v}}$, $b$ be an approximate solution for the graph coloring problem of $\mathsf{ConstructGraph}({\bm{v}})$, and let $h^{\prime}=m_{\rm color}(b)$. Since $b$ is not optimal, we have $\Delta:=h^{\prime}-h\geq 0$, where $h$ is the chromatic number of the graph. Due to Remark 6, we see that $m_{\rm depth}(\mathsf{ToCircuit}(b))-m_{\rm depth}({\bm{v}}_{\rm opt})\leq\Delta$, meaning that any approximate solution $b$ of the graph coloring leads to a circuit with smaller or equal gap to the minimum depth.
@@ 4 Applications
In Section 3, we have not specified conditions for operations to be simultaneously executable since the parallelizability depends on the underlying hardware. To examine the applications however, the condition has to be fixed. We begin with the following assumption that two gates sharing no common qubit can be executed simultaneously:
Assumption 9
@@
$\sum_{x}\lvert x\rangle\lvert 0\rangle\mapsto\sum_{x}\lvert x\rangle\lvert f(x)\rangle\enspace.$
@@ 4.1 Finite field multiplication
Remark 10
Using graph coloring, Toffoli depth of the multiplication circuit is reduced to as small as $12n-$ 1, from $44n-$ 4.
Figure 1: Depth optimization of multiplication in $\mathbb{F}_{2^{n}}$ by solving graph coloring. Red dots are obtained by using $\mathsf{DSatur}$ algorithm for graph coloring, black dashed line is the lower bound, and blue solid line is Maslov et al’s original design.
Maslov et al. originally proposed a circuit for binary field multiplication, but it can easily be extended to any $p>2$. The binary field multiplication circuit takes as inputs three quantum registers holding $a\in\mathbb{F}_{2^{n}}$, $b\in\mathbb{F}_{2^{n}}$, and 00, respectively on $n$ qubits each, and outputs $a,b,a\cdot b$. It requires $n^{2}$ Toffoli gates and $O(n^{2})$ CNOT gates, with $44n-$ 4 Toffoli depth. An important feature of the design is that no extra qubit is involved throughout the operation other than the initial $3n$ qubits.
@@
${\bm{d}}=\left[\begin{array}[]{c}a_{0}b_{0}\\ a_{1}b_{0}+a_{0}b_{1}\\ a_{2}b_{0}+a_{1}b_{1}+a_{0}b_{2}\\ a_{3}b_{0}+a_{2}b_{1}+a_{1}b_{2}+a_{0}b_{3}\end{array}\right],~{}{\bm{e}}=\left[\begin{array}[]{c}a_{3}b_{1}+a_{2}b_{2}+a_{1}b_{3}\\ a_{3}b_{2}+a_{2}b_{3}\\ a_{3}b_{3}\end{array}\right],~{}Q=\left[\begin{array}[]{ccc}1&0&0\\ 1&1&0\\ 0&1&1\\ 0&0&1\\ \end{array}\right]\enspace,$
@@
$\lvert{\bm{a}}\rangle\lvert{\bm{b}}\rangle\lvert 0\rangle\stackrel{{\scriptstyle\text{\tiny Phase 1}}}{{\longmapsto}}\lvert{\bm{a}}\rangle\lvert{\bm{b}}\rangle\lvert{\bm{e}}\rangle\stackrel{{\scriptstyle\text{\tiny Phase 2}}}{{\longmapsto}}\lvert{\bm{a}}\rangle\lvert{\bm{b}}\rangle\lvert Q{\bm{e}}\rangle\stackrel{{\scriptstyle\text{\tiny Phase 3}}}{{\longmapsto}}\lvert{\bm{a}}\rangle\lvert{\bm{b}}\rangle\lvert{\bm{d}}+Q{\bm{e}}\rangle\enspace.$
Figure 2 illustrates the procedures. Focusing on the Toffoli depth, in $\mathbb{F}_{2^{n}}$, the authors estimated the depth to be $32n-$ 3, 00, $12n-$ 1 in Phase 1, Phase 2, and Phase 3, respectively, in total $44n-$ 4.
Now, notice that all the additions occurring in Phase 1 and Phase 3 are commuting. The graph coloring strategy is applied to Phase 1 and Phase 3 in $\mathbb{F}_{2^{n}}$ for $n\leq 512$, adopting the solver $\mathsf{DSatur}$ which is a polynomial time heuristic algorithm for graph coloring [25]. Figure 1 summarizes the result. (Plots for $n>64$ are not shown in the figure, but all of them do not deviate at all from $12n-$ 1.) For $n\geq 6$, a lower bound on the Toffoli depth of $12n-$ 1 can be deduced from an elementary argument, [^1] It is the sum of the lengths of the bases of two triangles in Figure 2. For $n<6$, the lower bound is larger than $12n-$ 1 since a right triangle of base length $\leq 4$ (which appears for $n\leq 5$) cannot be fully parallelized. and $\mathsf{DSatur}$ has found circuits meeting the lower bound for most $n$ tested. [^2] For some small $n\;(<20)$ the corresponding graphs are rather denser in edge density, and the adopted $\mathsf{DSatur}$ code [33] has failed to meet the bound.
@@
As a side note, although not presented in this work, we have developed a way to build a circuit for the multiplication with the Toffoli depth of $2n$, but one for $12n-$ 1 has not been successful.

### v2

@@ Abstract
This work shows that minimizing the depth of a quantum circuit composed of commuting operations reduces to a vertex coloring problem on an appropriately constructed graph, where gates correspond to vertices and edges encode non-parallelizability. The reduction leads to algorithms for circuit optimization by adopting any vertex coloring solver as an optimization backend. The approach is validated by numerical experiments as well as applications to known quantum circuits, including finite field multiplication and QFT-based addition.
@@ 1 Introduction
Quantum circuits composed of finitely many unitary operations that mutually commute arise in several computational settings, including instantaneous quantum polynomial [1, 2], commuting local Hamiltonians [3, 4], and oracle constructions in quantum algorithms [5, 6]. Throughout this work, we refer to such circuits as commuting circuits. Because the underlying unitary operations mutually commute, such circuits are naturally amenable to parallelization, i.e., certain operations can be executed in parallel whenever the hardware supports it. Hardware-dependent constraints therefore determine which operations can be executed concurrently, and consequently lead to the notion of depth of a circuit.
In circuit synthesis, a function can generally be implemented by multiple constructions, and selecting one that minimizes resource usage is a central consideration in circuit design [7]. Under a fixed number of qubits, a design with smaller depth is typically preferred, as it reflects better utilization of the available parallelism. For this reason, among others, optimizing the depth of a circuit is an active area of research [7, 8, 9, 10, 11].
This study establishes a formal connection between depth optimization of commuting circuits and a well-known NP-hard problem in graph theory. The results are summarized as follows:
@@
    Two concrete algorithms are numerically tested to validate the method.
  • $\bullet$
    The applicability is demonstrated by optimizing circuits for finite field multiplications and additions by quantum Fourier transform (QFT). The first application exhibits the improvement upon Maslov et al.’s design [5] and the second one gives tradeoff options for Draper’s quantum addition [6].
The paper is organized as follows. Section 2 introduces the necessary preliminaries. Section 3 presents the main reduction and its proofs, followed by algorithms in Section 4. Section 5 discusses applications of the method, and Section 6 concludes the paper.
@@ 2.2 Gate Ordering Problem
Definition 1 (Commuting operations).
@@
Definition 2 (Depth of a circuit).
@@
Definition 3 (Gate ordering).
@@ 2.3 Graph coloring
Definition 4 (Proper coloring).
@@
Definition 5 (Vertex coloring problem).
Given a graph $G=(V,E)$, a vertex coloring problem is to find a colored graph $(V,E,c)\in\mathbf{Col}(G)$ such that $|\{c(v):v\in V\}|$ is minimized.
@@ 2.4 Related works
Circuit optimization is one of the active research areas in the quantum computing community. On the one hand, it is to ask theoretically how small a circuit can be, but on the other hand, finding better optimization is closely related to overcoming or mitigating noise problems [22]. The aspect of optimization varies, for example $T$-depth optimization of Clifford+$T$ circuits [23], pattern matching [24], local synthesis [25], or design automation [26]. This work pursues an optimization by the reordering of gates, in which the goal and the approach are different from previous works. The closest notion in the literature is the commutation analysis (built-in function of Qiskit [27]) whose purpose is to find cancellations or mergers, or to lower the overhead accounting for a hardware topology. We do not consider cancellations or so, but are interested in how much parallelization is achievable solely by reordering.
There exist works on quantum circuit optimization utilizing coloring in graph theory, by looking qubits as vertices. For example, edge coloring of a graph has been used to optimize the depth of Clifford circuits [28], or binary field inversion [29]. Such an approach is natural in the sense that once wires (qubits) are viewed as vertices and gates are considered as edges interconnecting the wires, the result of edge coloring tells which gates are parallelizable because two edges being connected to the same vertex implies they share the same qubit as the reference. Moreover, in cases where gates operating on more than two qubits are used, the approach can be generalized by considering hypergraphs where an edge can connect more than two vertices. For example in [30], the lower bound on the depth of the single optimization layer is given in terms of the chromatic index of the hypergraph [31].
Graph coloring has been used to optimally group Pauli strings to minimize the number of measurements in variational quantum eigensolver [32, 33, 34, 35, 36]. Given a set of Pauli strings, one tries to partition them into subsets such that all strings within each subset commute, while minimizing the number of subsets. The strings in each subset are simultaneously diagonalizable, allowing simultaneous measurement. This problem is reduced to solving graph coloring where the edges are drawn if two strings (vertices) are not commuting. The optimal partitioning of a set of Pauli strings is also considered in reducing the depth of a circuit for the time evolution of Sachdev-Ye-Kitaev model [37]. The authors successfully reduce the depth of the circuit by using the graph coloring algorithm.
We extend the idea by considering a circuit consisting only of commuting operations. The optimization approach developed in this paper uses an algorithm for graph coloring as the core routine. Any graph coloring algorithm can be used, for example including ones from previous works [38, 39, 40, 41], and it will soon be shown that the output of the graph coloring is directly related with the optimized circuit.
@@ 3 Reduction from gate ordering to graph coloring
The gate ordering problem is reduced to the graph coloring problem in this section. The opposite direction is also given in Appendix A, showing that the two problems are equivalent. Since the graph coloring problem is in NP-hard [14], it implies optimizing a commuting circuit is also hard.
@@ 3.1.2 Colored graph to circuit
Case (b) is straightforward, thus we leave a brief comment on the smaller case (a). Given a colored graph, unconnected vertices correspond to parallelizable gates, but the colors on these vertices can be different. Therefore, despite the different colors on the vertices, the corresponding gates from the output may happen to be in the same layer. An example is given in Appendix B.
@@ 3.2 Reduction
Recall that $G_{\cal V}$ is the graph drawn by Algorithm 1 for given $\bm{v}$}. Let $B_{\cal V}=\mathbf{Col}(G_{\cal V})$, which is the set of all colored graphs of $G_{\cal V}$.
Claim 7.
@@
$m_{\rm color}(b)=\min_{b^{\prime}\in B_{\cal V}}\big(m_{\rm color}(b^{\prime})\big)\kern 5.0pt.$ (1)
@@
$m_{\rm color}(g({\bm{a}}))=m_{\rm depth}({\bm{a}})<m_{\rm depth}(f(b))\leq m_{\rm color}(b)\kern 5.0pt,$
which contradicts to Eq. (1). ∎
@@
[…]
@@
17:return $(V,E,c)$ $\triangleright$ $c(v_{i})=c_{i}$
Let $\bm{v}$} be an input to $\mathsf{ToColoredGraph}$. Recall that the vertices are gates. Algorithm 3 sequentially identifies the parallelizable gates (vertices), and those vertices are identically colored in Line 7. Whenever a new color is introduced in Line 10, $d$ is also increased by 1. Since $d$ is increased only when $v_{j}$ is non-parallelizable with any of $v_{i},\ldots,v_{j-1}$ (deducible from Line 6), the final value of $d$ at the end is equal to the depth of the circuit. Therefore, we see that $\forall{\bm{a}}\in A_{\cal V}$, $m_{\rm depth}({\bm{a}})=m_{\rm color}(\mathsf{ToColoredGraph}({\bm{a}}))$. Time complexity of Algorithm 3 is $O(n^{2})$.
We have shown that Claim 7 does hold, thereby completing the reduction from gate ordering to graph coloring. A reduction for the opposite direction, from graph coloring to gate ordering, is less important as the main purpose of this work is to suggest a strategy for circuit optimization. Therefore we leave the opposite reduction in Appendix A.
@@ 3.3 Circuit optimization strategy
It can be deduced from inspecting the reduction steps that an approximate solution of graph coloring leads to at least as good approximations to the gate ordering problem. Let $\bm{v}$} be a commuting circuit, ${\bm{v}}_{\rm opt}$ be the optimal circuit of $\bm{v}$}, $b$ be an approximate solution for the graph coloring problem of $\mathsf{ConstructGraph}({\bm{v}})$, and let $h^{\prime}=m_{\rm color}(b)$. Since $b$ is not optimal, we have $\Delta:=h^{\prime}-h\geq 0$, where $h$ is the chromatic number of the graph. Due to Remark 6, we see that $m_{\rm depth}(\mathsf{ToCircuit}(b))-m_{\rm depth}({\bm{v}}_{\rm opt})\leq\Delta$, meaning that any approximate solution $b$ of the graph coloring leads to a circuit with smaller or equal gap to the minimum depth. Note however that the above argument holds only for approximate solutions. If the coloring is optimal, then the depth of the corresponding circuit is equal to the number of colors.
@@
4 Validation
In this section, the validity of the reduction strategy is put to the test by implementing Algorithm 4. Any coloring algorithm can be used, and we choose two well-known algorithms (solvers) and evaluate their performance.
Algorithm 4 $\mathsf{OptimizeCircuit}$
[…]
3:
[…]
7:return $\bm{v}$$^{\prime}$
4.1 Coloring solvers
4.1.1 DSatur
DSatur is a saturation-degree-based greedy algorithm [39]. Saturation degree of a vertex is the number of different colors used by its neighbors. Roughly described, at each step of the algorithm, an uncolored vertex with the highest degree is selected and assigned a color such that the use of a new color is suppressed if possible.
The time complexity is easily estimated. Let the number of vertices of an input graph $G$ be $n$. At each step, for a selected vertex, the algorithm inspects only its neighbors and then colors it. It repeats for all $n$ vertices, thus the running time is bounded by $O(n^{2})$.
About the number of colors of the output, let $\chi_{G}$ be the chromatic number of $G$. The algorithm is heuristic, and there is no general theoretical bound that characterizes how close DSatur’s output is to $\chi_{G}$. However, the empirical performance of DSatur has been studied extensively, and the benchmark results can be found in various literature [42, 43].
4.1.2 Backtracking
Backtracking is an algorithm for $k$-coloring, which asks if a given graph is colorable with at most $k$ different colors [44]. Basically, Backtracking performs a depth-first search over color assignments. It colors vertices one by one while respecting constraints. When a vertex cannot be assigned a color anymore without violating the constraints, the algorithm backtracks to change earlier color choices, effectively pruning that branch of the search space. The worst-case time complexity is known to be $O(k^{n})$ for $k$-coloring and it can be used to find the chromatic number by trying different $k$ at most $n$ times.
4.2 Lower bound on the depth of a commuting circuit
In Section 3, we have not specified conditions for operations to be simultaneously executable since the parallelizability depends on the underlying hardware. To analyze the algorithm however, the condition has to be fixed. We begin with the following assumption that two gates sharing no common qubit can be executed simultaneously:
Assumption 9.
When a commuting circuit is generated randomly (each gate is chosen randomly), the corresponding graph by Algorithm 1 becomes a random intersection graph[45]. We skip the details on the model, but take the bound on the chromatic number. Let $m$ be the number of qubits of a random commuting circuit and $G$ be the corresponding graph, then
$\chi(G)\geq\max_{1\leq q\leq m}|L_{q}|\kern 5.0pt,$ (2)
where $\chi(G)$ is the chromatic number of $G$, and $L_{q}$ is the set of gates that act on the $q$-th qubit. For example for a circuit ${\bm{v}}_{2}=(C\!X_{134},C\!X_{235},C\!X_{167},C\!X_{268})$, we have $L_{1}=\{C\!X_{134},C\!X_{167}\}$ and $L_{2}=\{C\!X_{235},C\!X_{268}\}$. Briefly explained, the largest set $L_{q}$ naturally forms the maximum clique [14] whose size provides the non-tight lower bound on the chromatic number.
It is worth noting that for certain parameter regimes- the number of gates, the number of qubits, the sizes of gates- the bound is nearly tight and polynomial time algorithms can meet the lower bound [46, 47, 48]. However, delving into the optimal solvability of commuting circuits is beyond the scope of this work.
4.3 Experiments
We generate commuting circuits with $m\in\{10,20,30\}$ qubits and $n\in\{10,20,...,200\}$ multi-controlled $Z$ gates. When generating a gate, we randomly choose at least three to at most $\lfloor m/2\rfloor$ qubits the gate will act on. The generated circuit is then fed to Algorithm 4 as an input. For each $m$ and $n$, 100 circuits are generated and optimized by the algorithm. The results are shown in Figure 1.
Figure 1: Optimization by coloring for randomly generated commuting circuits with $m\in\{10,20,30\}$ qubits and $n\in\{10,20,...,200\}$ gates. Blue, orange, green, and red markers denote, respectively, the averaged depths of the input circuit, the output circuit produced by DSatur, the output circuit produced by Backtracking, and the non-tight lower bound by Eq. (2).
In Figure 1, each data point represents an average over 100 independent trials. The blue marker indicates the depth of the input circuits, the orange (green) marker represents the depth of the output circuit obtained using the DSatur (Backtracking) solver, and the red marker denotes the non-tight lower bound given by Eq. (2). Since the running time of the exact solver is exponential in the number of gates, the green markers are present up to $n=50$. For all $(n,m)$ values where green markers are present, the orange markers lie unnoticeably slightly above or exactly on the green ones. How much more the two solvers deviate for larger circuit sizes is open for further experiment. Each set of circuit parameters ($n$, $m$, and gate sizes) may correspond to a distinct hardness regime for graph coloring in random intersection graphs [48]. Based on the observations, we hypothesize that gate ordering is easier for small numbers of qubits, as the polynomial time heuristic algorithm produced circuits that were nearly optimal. For larger numbers of qubits however, we cannot immediately draw the conclusion that the problems are harder since the lower bound is non-tight.
In addition, we conducted an experiment on the coloring algorithms within the Qiskit framework. Direct comparison is limited because Qiskit transpiler does not work as a gate ordering solver. Instead, we let Qiskit to decompose a commuting circuit, with and without the coloring algorithms applied, into Clifford+$T$ circuit and then let it optimize to reduce the number of $T$ gates. Details are given in Appendix C.
@@ 5 Applications
$\sum_{x}\lvert x\rangle\lvert 0\rangle\mapsto\sum_{x}\lvert x\rangle\lvert f(x)\rangle\kern 5.0pt.$
@@ 5.1 Finite field multiplication
Figure 2: Depth optimization of multiplication in $\mathbb{F}_{2^{n}}$ by solving graph coloring. Red markers are obtained by using DSatur algorithm for graph coloring, black dashed line is the lower bound, and blue solid line is Maslov et al’s original design.
Remark 10.
Using graph coloring, Toffoli-depth of the multiplication circuit is reduced to as small as 2n12n-1, from 4n44n-4.
Maslov et al. originally proposed a circuit for binary field multiplication, but it can easily be extended to any $p>2$. The binary field multiplication circuit takes as inputs three quantum registers holding $a\in\mathbb{F}_{2^{n}}$, $b\in\mathbb{F}_{2^{n}}$, and 0, respectively on $n$ qubits each, and outputs $a,b,a\cdot b$. It requires $n^{2}$ Toffoli gates and $O(n^{2})$ CNOT gates, with 4n44n-4 Toffoli-depth. An important feature of the design is that no extra qubit is involved throughout the operation other than the initial $3n$ qubits.
@@
${\bm{d}}=\left[\begin{array}[]{c}a_{0}b_{0}\\ a_{1}b_{0}+a_{0}b_{1}\\ a_{2}b_{0}+a_{1}b_{1}+a_{0}b_{2}\\ a_{3}b_{0}+a_{2}b_{1}+a_{1}b_{2}+a_{0}b_{3}\end{array}\right],\penalty 10000\ {\bm{e}}=\left[\begin{array}[]{c}a_{3}b_{1}+a_{2}b_{2}+a_{1}b_{3}\\ a_{3}b_{2}+a_{2}b_{3}\\ a_{3}b_{3}\end{array}\right],\penalty 10000\ Q=\left[\begin{array}[]{ccc}1&0&0\\ 1&1&0\\ 0&1&1\\ 0&0&1\\ \end{array}\right]\kern 5.0pt,$
@@
$\lvert{\bm{a}}\rangle\lvert{\bm{b}}\rangle\lvert 0\rangle\stackrel{{\scriptstyle\text{\tiny Phase 1}}}{{\longmapsto}}\lvert{\bm{a}}\rangle\lvert{\bm{b}}\rangle\lvert{\bm{e}}\rangle\stackrel{{\scriptstyle\text{\tiny Phase 2}}}{{\longmapsto}}\lvert{\bm{a}}\rangle\lvert{\bm{b}}\rangle\lvert Q{\bm{e}}\rangle\stackrel{{\scriptstyle\text{\tiny Phase 3}}}{{\longmapsto}}\lvert{\bm{a}}\rangle\lvert{\bm{b}}\rangle\lvert{\bm{d}}+Q{\bm{e}}\rangle\kern 5.0pt.$
Figure 3 illustrates the procedures. Focusing on the Toffol-depth, in $\mathbb{F}_{2^{n}}$, the authors estimated the depth to be 2n32n-3, 0, 2n12n-1 in Phase 1, Phase 2, and Phase 3, respectively, in total 4n44n-4.
Now notice that all the additions occurring in Phase 1 are commuting. Similarly, all the additions in Phase 3 are also commuting. The graph coloring strategy is applied to each Phase 1 and Phase 3 in $\mathbb{F}_{2^{n}}$ for $n\leq 512$, adopting the solver DSatur which is a polynomial time heuristic algorithm for graph coloring [39]. Total Toffoli-depth is the sum of the coloring results in two phases. Figure 2 summarizes the result. (Plots for $n>64$ are not shown in the figure, but all of them do not deviate at all from 2n12n-1.) For $n\geq 6$, a lower bound on the Toffoli-depth of 2n12n-1 can be deduced from an elementary argument, [^1] It is the sum of the lengths of the bases of two triangles in Figure 3. For $n<6$, the lower bound is larger than 2n12n-1 since a right triangle of base length $\leq 4$ (which appears for $n\leq 5$) cannot be fully parallelized. and DSatur has found circuits meeting the lower bound for most $n$ tested. [^2] For some small n(<20)n\;(<20) the corresponding graphs are rather denser in edge density, and the adopted DSatur code [54] has failed to meet the bound.